OpenStudy (anonymous):

Solve. c minus 9 is greater than twelve c is greater than 3 c is greater than 9 c is greater than twelve c is greater than twenty-one

OpenStudy (anonymous):

hey just try to convert the wordings into mathematics notation you ll obtain the answer...

OpenStudy (anonymous):

so c is greater than 9

OpenStudy (anonymous):

C-9= 12 +9 +9 --------- C = 3 ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nope c is the value where when you subtract 9 from it , the value you obtained is greater than 12